1) If four particles have velocities +1. -4, 5, 10 m/s respectively, what is the average speed?

2) How much ice at 0 degrees Celsius must be added to a container of 200 grams of water at 80 degrees Celsius to reduce the temperature to 45 degrees Celsius?

3) A real engine takes in heat at 966, and discharges heat at 300 K. We know its efficiency must be less than?
Answer #1

1)

here , let the average speed is Vavg

Vavg = total sum of energy/total number of particle

Vavg = ( 1 - 4 + 5 + 10)/4

Vavg = 3 m/s

the average velocity is 3 m/s

2)

let the mass of ice needed is m

energy released water = energy gained by ice

m * Lv + m * Swater * del T = mass of water * Swater * del T

m * (334 + 4.186 * (45 - 0 )) = 200 * 4.186 * (80 - 45)

solving for m

m = 56 gm

the mass of ice needed is 56 gm

3)

Here , Th = 966 K

Tc = 300 K

efficiency = 1 - Tc/Th

efficiency = 1- 300/966

efficiency = 0.689

the efficiency must be less than 0.689